OpenStudy (anonymous):

solve for x --- arcsin(2x)^1/2 = arccos(x)^1/2

OpenStudy (anonymous):

Nobody?

OpenStudy (anonymous):

im getting x is: \[\frac{1}{\sqrt{5}}\], but i need to verify my answer. one sec.

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=solve+arcsin(2x)^1%2F2+%3D+arccos(x)^1%2F2 This is the solution. I'll let joe help you do the analytic part.

OpenStudy (anonymous):

You're right joe^^^^

OpenStudy (anonymous):

book says x=1/3......

OpenStudy (anonymous):

Then you have type the problem wrong :P

OpenStudy (anonymous):

is it this? \[(\sin^{-1}(2x))^{\frac{1}{2}} = (\cos^{-1}(x))^{\frac{1}{2}}\]

OpenStudy (anonymous):

cause thats what i solved.

OpenStudy (anonymous):

arcsin and sin-1 are the same yes?

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

note, what i typed is different than: \[\sin^{-1}((2x)^\frac{1}{2}) = \cos^{-1}((x)^\frac{1}{2})\]

OpenStudy (anonymous):

arcsin\[\sqrt{2x}\]+arcos\[\sqrt{x}\] solve for x

OpenStudy (anonymous):

ah, its the second one then.

OpenStudy (anonymous):

I don't know why it's put in a column like that...

OpenStudy (anonymous):

It automatically spaces it that way.

OpenStudy (anonymous):

its like that, but i get the problem now.

OpenStudy (anonymous):

so the way i solved it was i said let:\[y = \sin^{-1}(\sqrt{2x}), y = \cos^{-1}(\sqrt{x})\] we can set them both equal to y since they both equal each other. Taking the sine of both sides of the first equation, and cosine of both sides in the second gives: \[\sin(y) = \sqrt{2x}, \cos(y) = \sqrt{x} \iff \sin^2(y) = 2x, \cos^2(y) = x\]Now using the trig identity:\[\sin^2(y)+\cos^2(y) = 1\]we obtain: \[2x+x = 1 \iff 3x = 1 \iff x = \frac{1}{3}\]

OpenStudy (anonymous):

steps are unavailable on the wolfram website.....

OpenStudy (anonymous):

very good. thank you. it makes prefect sense now